Find the Cartesian equations of the lines or curves whose polar equations are (a) r = a (b) r = a cos 2θ
(c) r2 = a2 sin 2θ (e) r = a sec(θ − α)
1
Expert's answer
2020-04-21T15:40:08-0400

Note: "r=\sqrt{x^2+y^2}; sin(\theta)=\frac{y}{\sqrt{x^2+y^2}}; cos(\theta)=\frac{x}{\sqrt{x^2+y^2}}"

a) "r=a"

"\sqrt{x^2+y^2}=a\\
x^2+y^2=a^2"


b)"r=acos(2\theta)"

"r=a(2cos^2(\theta)-1)\\
\sqrt{x^2+y^2}=a(2(\frac{x}{\sqrt{x^2+y^2}})^2-1)\\
\sqrt{x^2+y^2}=a(\frac{2x^2}{x^2+y^2}-1)\\
\sqrt{x^2+y^2}=a\frac{x^2-y^2}{x^2+y^2}"

"x^2+y^2=a^2\frac{(x^2-y^2)^2}{(x^2+y^2)^2}\\
(x^2+y^2)^3=a^2(x^2-y^2)^2"


c) "r^2=a^2sin(2\theta)"

"r^2=2a^2sin(\theta)cos(\theta)\\
x^2+y^2=2a^2\frac{y}{\sqrt{x^2+y^2}}\frac{x}{\sqrt{x^2+y^2}}\\
x^2+y^2=2a^2\frac{xy}{x^2+y^2}\\
(x^2+y^2)^2=2a^2xy"


e)"r=asec(\theta-\alpha)"

"r=\frac{a}{cos(\theta-\alpha)}\\
rcos(\theta-\alpha)=a"

"r(cos(\theta)cos(\alpha)+sin(\theta)sin(\alpha))=a"

"\sqrt{x^2+y^2}(\frac{x}{\sqrt{x^2+y^2}}cos(\alpha)+\frac{y}{\sqrt{x^2+y^2}}sin(\alpha))=a\\
xcos(\alpha)+ysin(\alpha)=a\\"